OpenStudy (anonymous):

* Advance Algebra With Financial Application * Kate bought a washer from a rental center for $875.63. She makes 12 monthly payments of $88.38 with her credit card. The rental center charges $2.50 for every purchase made with a credit card. She also had to pay late fees in the amount of $45 four different times. What is the total cost of the washer? $1,090.56 $1,060.56 $1,270.56 $1,240.56

OpenStudy (anonymous):

Been trying to figure it out and I cant get nothing :c

OpenStudy (nincompoop):

well… it's easy if you can follow the addition

OpenStudy (nincompoop):

let us try this the crude way

OpenStudy (nincompoop):

we can tell that in the end Kate paid more than the price tag of the washer. the sale cost is 875.63 but she pays 88.38 this means that in twelve payments, she would have paid 1060.56 then she pays extra 2.50 for every payment with her card. multiplying this by 12 gives 30 summing this, she's already looking at 1060.56 + 30 = 1090.56 but since Kate is delinquent 4 times with her payment and get late payment fee of 45. multiplying this by 4 gives 1090.56 + 180 = ? let us run the totals (88.38 x 12) + (2.50 x 12) + (45.00 x 4) =

OpenStudy (nincompoop):

so far, there's nothing advanced in the problem. they're just mere simple accounting